# Vision Mamba in Remote Sensing: A Comprehensive Survey of Techniques, Applications and Outlook

### v1

@@ Abstract
Deep learning has profoundly transformed remote sensing, yet prevailing architectures like Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) remain constrained by critical trade-offs: CNNs suffer from limited receptive fields, while ViTs grapple with quadratic computational complexity, hindering their scalability for high-resolution remote sensing data. State Space Models (SSMs), particularly the recently proposed Mamba architecture, have emerged as a paradigm-shifting solution, combining linear computational scaling with global context modeling. This survey presents a comprehensive review of Mamba-based methodologies in remote sensing, systematically analyzing about 120 studies to construct a holistic taxonomy of innovations and applications. Our contributions are structured across five dimensions: (i) foundational principles of vision Mamba architectures, (ii) micro-architectural advancements such as adaptive scan strategies and hybrid SSM formulations, (iii) macro-architectural integrations, including CNN-Transformer-Mamba hybrids and frequency-domain adaptations, (iv) rigorous benchmarking against state-of-the-art methods in multiple application tasks, such as object detection, semantic segmentation, change detection, etc. and (v) critical analysis of unresolved challenges with actionable future directions. By bridging the gap between SSM theory and remote sensing practice, this survey establishes Mamba as a transformative framework for remote sensing analysis. To our knowledge, this paper is the first systematic review of Mamba architectures in remote sensing. Our work provides a structured foundation for advancing research in remote sensing systems through SSM-based methods. We curate an open-source repository (https://github.com/BaoBao0926/Awesome-Mamba-in-Remote-Sensing) to foster community-driven advancements.
@@ I Introduction
To address these challenges, state space models (SSMs)-based methods have emerged as a promising alternative, offering both linear computational complexity and global modeling capabilities. These two characteristics of SSMs precisely overcome the limitations of CNNs and ViTs in remote sensing applications. Rooted in classical system theory, SSMs have found widespread applications across disciplines including reinforcement learning [8], computational neuroscience [9], and control systems [10]. The SSM maps input sequences to latent states that encapsulate historical context, enabling sequential prediction according to the hidden state. Some work has tried to incorporate SSM into the deep learning framework [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Early SSM implementations faced computational bottlenecks until Structured State Space sequence (S4) model [11] addressed these limitations through parameterized state matrices. Then, Mamba [20] introduces Selective State Space for Sequences (S6) that incorporates dynamic time-aware mechanisms via regressed step size parameters, enabling context-aware information propagation or forgetting through hidden states. The Mamba architecture [20] further advanced this paradigm through simplified gated SSM blocks, achieving state-of-the-art (SOTA) performance. With its linear scalability for high-resolution imagery, ability to capture long-range spatial dependencies and exceptional capability of feature representation, Mamba-based architectures have demonstrated strong potential as a next-generation solution for remote sensing tasks, bridging the gap between computational efficiency and global modeling.
@@
  • Scan Strategy: We propose a novel taxonomy for scan strategies (crucial for processing 2D/3D imagery as 1D sequences) comprising five elements: preprocessing, scan sampling, scan direction, scan pattern, and postprocessing. This framework provides the first comprehensive treatment that includes preprocessing and postprocessing stages into scan strategy.
@@
  • Substitution in Framework Adaptation: We survey the incorporation of Mamba blocks into established frameworks including U-Net [42], YOLO [43] and Diffusion models [44].
  • Learning Paradigms: In addition to the conventional supervised learning, our review encompasses applications of unsupervised, self-supervised, and prompt learning paradigm.
@@ II-A State Space Model
This linear recurrence calculation can be accelerated by SMM convolution kernel as follows:
@@ II-B Mamba: Selective SSM
The core building block of Mamba is a simplified attention architecture. Unlike conventional SSM designs that employ a stacked combination of linear attention-like blocks and MLP blocks, similar to Transformer [4] architectures, Mamba unifies these two fundamental components into a single structure, forming the Mamba block. As depicted in Fig. 2a, the Mamba block can be analyzed from two distinct perspectives. First, it substitutes the multiplicative gating mechanism in linear attention-like or H3 [14] blocks with an activation function. Second, it integrates the SSM transformation directly into the primary computational pathway of the MLP block. The overall architecture of Mamba is composed of multiple Mamba blocks, interspersed with standard normalization layers and residual connections. In detail, the activation function uses SiLU [45] and Swish activation [46], and we normally call the activation score in the gating pathway as $z$. This architecture effectively captures long-range dependencies while maintaining the linear scalability characteristic of SSMs with respect to sequence length, becoming a new promising foundation model for CV.
(a) The architecture of Mamba, a simplified attention block that integrates the SSM transformation directly into the primary computational pathway of the MLP block and directly uses activation function on the gating pathway.
(b) Mamba-2 Block Architectures: (Left) Sequential block that operations are applied in series. (Right) Parallel blocks that selected operations executes concurrently before fusion.
@@ II-C1 Architectural Innovations
Key advancements include: 1) Parallel Parameterization: Co-projecting SSM parameters ($A,B,C$) and values ($X$) akin to attention’s $QKV$ mappings, reducing synchronization bottlenecks by 50% compared to Mamba-1 [20]; 2) Multi-Input SSM (MIS) Heads: Sharing $B/C$ projections across channels while keeping $A$ head-specific, enabling 8$\times$ larger state dimensions (e.g., $N=512$) with minimal overhead; 3) Hybrid Blocks: As illustrated in Fig. 2b, Mamba-2 supports both sequential and parallel configurations of its core components (e.g., convolutional layers, SSM transformations, and nonlinear activations), allowing flexible trade-offs between computational efficiency and model capacity.
@@ II-C2 Efficiency and Performance
Mamba-2 achieves Pareto dominance over Transformers and Mamba-1, attaining 6.09 perplexity on the Pile (2.7B parameters) with 3$\times$ faster training than FlashAttention-2 at 16K context [48, 49]. Its blockwise SSD algorithm splits sequences into chunks, combining intra-chunk matrix multiplications and inter-chunk state transitions for 2–8$\times$ speedups. System optimizations include tensor parallelism with single all-reduce per block and sequence-parallel state propagation for billion-token sequences. Evaluations highlight 98% accuracy on memory-intensive tasks (e.g., multi-query associative recall), outperforming Mamba-1 by 16% [50].
@@ III Vision Mamba Backbones
Considering one fundamental challenge in adapting SSMs to vision tasks lies in effectively transforming 2D spatial relationships into sequential representations, numerous work has focused on modifying scan strategies to increase the model’s understanding of non-causal 2D data. Building upon the approaches of Vim and VMamba, some work focus on scan strategies (discussed in Section IV-B4). Mamba-ND [51] generalizes scan strategies to accommodate 3D data by employing block-level alternation of scan directions along the height, width, and time axes. LocalVMamba [52] integrates local window-based scanning (Fig. 8.D) with learnable pathway weights through DARTS-inspired architecture search [53]. PlainMamba [54] and FractalMamba [55] employ a continuous 2D scan pattern (Fig. 8.B) and Hilbert curve pattern (Fig. 8.E), respectively, aiming to enhance contextual understanding capabilities and address the issue of discontinuity. Subsequently, some researchers have leveraged preprocessing methods (discussed in Section IV-C2) to enhance the performance of SSMs. MS-VMamba [56] applies multi-scale SSMs, while GroupMamba [57] divides the input feature into four groups in channel dimensions and applies a separate SSM independently to each group.
@@
Some researchers have developed Mamba-based architecture on designs that are proven effective in CNN/Transformer blocks. MambaMixer [58] follows the core idea of MLP-Mixer [59] to combine Mamba-based token mixing with channel mixing, and follows the design of DenseNet [60] and DenseFormer [61] to allow blocks to access the previous features. Following [62], MambaR [63] evenly inserts register tokens into the input sequences and recycle registers for final decision predictions, which can help model focus on more semantic regions. ARM [64] follows the GPT families [65, 66, 67], which utilizes the autoregressive modeling self-supervised framework. Similar to EViT [68], MambaPruning [69] introduces a token pruning method specifically designed for Mamba-based model. Inspired by [70], ShuffleMamba [71] introduces a training regularization technique that randomly shuffles token sequences during training to improve positional transformation invariance and overall model performance.
@@
On one hand, some papers focus on the micro-architecture designs to enhance model’s ability. VSSD [72] introduces non-causal processing for image data by decoupling interaction magnitudes between hidden states and tokens. SparX-Mamba [73] introduces a efficient sparse cross-layer feature aggregation method specifically designed for SSMs. Vim-F [74] integrates frequency domain information into the basic block to achieve a better global receptive field. On the other hand, some work focuses on the macro-architecture advancement. SiMBA [75] utilizes hybrid sequence-channel modeling, effectively resolving the instability issues in large-size vision Mamba. StableMamba [76] establishes a robust interleaved Mamba-Attention framework that successfully scales SSMs to over 100 million parameters without resorting to knowledge distillation techniques. MambaVision [77] redesigns the Mamba architecture and integrates CNNs and ViT block on the architecture level, aiming at enhancing modeling capability.
@@ IV Micro-Architecture Advancement
This section focus on the inner mechanism within micro-architecture of Mamba block. Three core points are summarized: (1) SSM Formula Improvement (Section IV-A); (2) Scan Strategy (Section IV-B) and (3) Multi-Modal and Bi-Temporal Feature Interaction via Mamba (Section IV-C).
@@ IV-A SSM Formula Improvement
Figure 2: Overview of Recent Advancements in SSM Formulations. This figure categorizes improvements into four major aspects: (a) Context-SSM [78, 79] for local feature enhancement, integrating local context through CNNs; (b) Cross-SSM [80, 35, 36, 81, 82] for multi-modal and bi-temporal feature interaction, enabling cross-modal/temporal feature interaction at the parameter level; (c) New discretization methods for SSM [83], introducing second-order Runge-Kutta (RK2) for improved continuous-time approximation; and (d) Alternative hidden state definitions [84], utilizing a Minimum Search Tree (MST) to redefine hidden state transitions. These refinements collectively enhance the efficiency and expressiveness of SSMs in various applications
@@ IV-A1 Context-SSM for Local Feature Enhancement
Context-SSM aims to naturally incorporate local context information into modeling long-range dependencies. Mamba-MOCO [78], building upon insights from MambaIRv2 [85], identifies two critical insights for improving SSMs in CV: (1) the output matrix $C$ in Eq. 7 functions as a ”Query” for local context information, which can address the inherent limitation of transforming 2D images into 1D sequences; (2) enhancing local context allows the model to better focus on local interactions while preserving its ability to model global dependencies. In response to these insights, Mamba-MOC [78] introduces the Context-SSM, as depicted in Fig. 2.a. Context-SSM incorporates two new parameters, $P$ and $D$, which are generated by multi-scale CNNs and a $1\times 1$ CNN. The parameter $P$ and $D$ capture local context information, ensuring that the model can effectively capture both local context and long-range dependencies for the input data. Similarly, AFA-Mamba [79] simply adds local context feature $eg(L_{p})$ with output equation together, which improves the ability of understanding local context.
@@ IV-B Scan Strategy
Scan Strategy Papers
@@
TABLE I: This table summarizes a total of 44 different scan strategies utilized in the field of remote sensing. All acronyms are clearly illustrated in Fig. 4-10. The terms WB and BB denote ’Within Block’ and ’Between Block’, respectively, and are discussed in detail in Section IV-B1. Arrows in the direction column indicate the starting token and initial scan directions, which are further explained in Section IV-B3.
@@
Figure 3: Illustration of the scan strategy pipeline, comprising five key components, including feature preprocessing, scan sampling, scan directions, scan pattern, and feature post-processing. For clarity, an additional example is provided demonstrating the case without any preprocessing or postprocessing operations.
@@ IV-B1 Feature Preprocessing
Unlike Transformers [4, 170] processing tokens in parallel, Mamba performs sequential computation, meaning that the initial scan order and the starting token influences subsequent token interactions. This category aims to optimize spatial-spectral feature extraction through feature map rearrangement [122, 150, 113]. HSRMamba [122] introduces the Global Spectral Reordering Mechanism (GSRM), which rearranges spectral features based on the average correlation matrix values between spectrum, ensuring that highly correlated pixels are positioned closer together, as illustrated in Fig. 4.E. This improves long-range spectral feature modeling. VMambaSCI [150] proposes Spectral Exchange (SE), which reorders spectral features using a predefined sequence, as shown in Fig. 4.F. This enhances the long-range modeling of spectral features within the same spatial channel. DTAM [113] introduces Dynamic Token Argumentation (DTA) to enhance global feature learning. As depicted in Fig. 4.G, DTA consists of two routes: the left route activates object-related tokens while masking the others, and the right route activates contextual tokens while using random shuffling and Adaptive Instance Normalization (AdaIN) [171] for augmentation. This approach helps the model focus on meaningful object-related features while mitigating spectral variability caused by acquisition conditions.

### v2

@@ Abstract
Deep learning has profoundly transformed remote sensing, yet prevailing architectures like Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) remain constrained by critical trade-offs: CNNs suffer from limited receptive fields, while ViTs grapple with quadratic computational complexity, hindering their scalability for high-resolution remote sensing data. State Space Models (SSMs), particularly the recently proposed Mamba architecture, have emerged as a paradigm-shifting solution, combining linear computational scaling with global context modeling. This survey presents a comprehensive review of Mamba-based methodologies in remote sensing, systematically analyzing about 120 Mamba-based remote sensing studies to construct a holistic taxonomy of innovations and applications. Our contributions are structured across five dimensions: (i) foundational principles of vision Mamba architectures, (ii) micro-architectural advancements such as adaptive scan strategies and hybrid SSM formulations, (iii) macro-architectural integrations, including CNN-Transformer-Mamba hybrids and frequency-domain adaptations, (iv) rigorous benchmarking against state-of-the-art methods in multiple application tasks, such as object detection, semantic segmentation, change detection, etc. and (v) critical analysis of unresolved challenges with actionable future directions. By bridging the gap between SSM theory and remote sensing practice, this survey establishes Mamba as a transformative framework for remote sensing analysis. To our knowledge, this paper is the first systematic review of Mamba architectures in remote sensing. Our work provides a structured foundation for advancing research in remote sensing systems through SSM-based methods. We curate an open-source repository (https://github.com/BaoBao0926/Awesome-Mamba-in-Remote-Sensing) to foster community-driven advancements.
@@ I Introduction
To address these challenges, state space models (SSMs)-based methods have emerged as a promising alternative, offering both linear computational complexity and global modeling capabilities. These two characteristics of SSMs precisely overcome the limitations of CNNs and ViTs in remote sensing applications. Rooted in classical system theory, SSMs have found widespread applications across disciplines including reinforcement learning [8], computational neuroscience [9], and control systems [10]. The SSM maps input sequences to latent states that encapsulate historical context, enabling sequential prediction according to the hidden state. Some work has tried to incorporate SSM into the deep learning framework [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Early SSM implementations faced computational bottlenecks until the Structured State Space sequence (S4) model [11] addressed these limitations through parameterized state matrices. Then, Mamba [20] introduces Selective State Space for Sequences (S6) that incorporates dynamic time-aware mechanisms via regressed step size parameters, enabling context-aware information propagation or forgetting through hidden states. The Mamba architecture [20] further advanced this paradigm through simplified gated SSM blocks, achieving state-of-the-art (SOTA) performance. With its linear scalability for high-resolution imagery, ability to capture long-range spatial dependencies and exceptional capability of feature representation, Mamba-based architectures have demonstrated strong potential as a next-generation solution for remote sensing tasks, bridging the gap between computational efficiency and global modeling.
@@
  • Scan Strategy: We propose a novel taxonomy for scan strategies (crucial for processing 2D/3D imagery as 1D sequences) comprising five elements: preprocessing, scan sampling, scan direction, scan pattern, and postprocessing. This framework provides the first comprehensive treatment that includes preprocessing and postprocessing stages into the scan strategy.
@@
  • Substitution in Framework Adaptation: We survey the incorporation of Mamba blocks into established frameworks, including U-Net [42], You Only Look Once (YOLO) [43] and Diffusion models [44].
  • Learning Paradigms: In addition to the conventional supervised learning, our review encompasses applications of unsupervised, self-supervised, and prompt learning paradigms.
@@ II-A State Space Model
This linear recurrence calculation can be accelerated by SSM convolution kernel as follows:
@@ II-B Mamba: Selective SSM
The core building block of Mamba is a simplified attention architecture. Unlike conventional SSM designs that employ a stacked combination of linear attention-like blocks and MLP blocks, similar to Transformer [4] architectures, Mamba unifies these two fundamental components into a single structure, forming the Mamba block. As depicted in Fig. 2.a, the Mamba block can be analyzed from two distinct perspectives. First, it substitutes the multiplicative gating mechanism in linear attention-like or H3 [14] blocks with an activation function. Second, it integrates the SSM transformation directly into the primary computational pathway of the MLP block. The overall architecture of Mamba is composed of multiple Mamba blocks, interspersed with standard normalization layers and residual connections. In detail, the activation function uses SiLU [45] and Swish activation [46], and we normally call the activation score in the gating pathway as $z$. This architecture effectively captures long-range dependencies while maintaining the linear scalability characteristic of SSMs with respect to sequence length, becoming a new promising foundation model for CV.
Figure 2: The architecture of Mamba-1 [20] and Mamba-2 [21]. (a) Mamba-1, a simplified attention block that integrates the SSM transformation directly into the primary computational pathway of the MLP block and uses an activation function on the gating pathway. (b) Sequential Mamba-2 Block, where operations are applied in series, and (c) Parallel Mamba-2 Block, where selected operations execute concurrently before fusion.
@@ II-C1 Architectural Innovations
Key advancements include: 1) Parallel Parameterization: Co-projecting SSM parameters ($A,B,C$) and values ($X$) akin to attention’s $QKV$ mappings, reducing synchronization bottlenecks by 50% compared to Mamba-1 [20]; 2) Multi-Input SSM (MIS) Heads: Sharing $B/C$ projections across channels while keeping $A$ head-specific, enabling 8$\times$ larger state dimensions (e.g., $N=512$) with minimal overhead; 3) Hybrid Blocks: As illustrated in Fig. 2.b and Fig. 2.c, Mamba-2 supports both sequential and parallel configurations of its core components (e.g., convolutional layers, SSM transformations, and nonlinear activations), allowing flexible trade-offs between computational efficiency and model capacity.
@@ II-C2 Efficiency and Performance
Mamba-2 achieves Pareto dominance over Transformers and Mamba-1, attaining 6.09 perplexity on the Pile (2.7B parameters) with 3$\times$ faster training than FlashAttention-2 at 16K context [48, 49]. Its blockwise SSD algorithm splits sequences into chunks, combining intra-chunk matrix multiplications and inter-chunk state transitions for 2–8$\times$ speedups. System optimizations include tensor parallelism with a single all-reduce per block and sequence-parallel state propagation for billion-token sequences. Evaluations highlight 98% accuracy on memory-intensive tasks (e.g., multi-query associative recall), outperforming Mamba-1 by 16% [50].
@@ III Vision Mamba Backbones
Considering one fundamental challenge in adapting SSMs to vision tasks lies in effectively transforming 2D spatial relationships into sequential representations, numerous works have focused on modifying scan strategies to increase the model’s understanding of non-causal 2D data. Building upon the approaches of Vim and VMamba, some work focuses on scan strategies (discussed in Section IV-B4). Mamba-ND [51] generalizes scan strategies to accommodate 3D data by employing block-level alternation of scan directions along the height, width, and time axes. LocalVMamba [52] integrates local window-based scanning (Fig. 9.D) with learnable pathway weights through DARTS-inspired architecture search [53]. PlainMamba [54] and FractalMamba [55] employ a continuous 2D scan pattern (Fig. 9.B) and Hilbert curve pattern (Fig. 9.E), respectively, aiming to enhance contextual understanding capabilities and address the issue of discontinuity. Subsequently, some researchers have leveraged preprocessing methods (discussed in Section IV-C2) to enhance the performance of SSMs. MS-VMamba [56] applies multi-scale SSMs, while GroupMamba [57] divides the input feature into four groups in channel dimensions and applies a separate SSM independently to each group.
@@
Some researchers have developed Mamba-based architecture on designs that are proven effective in CNN/Transformer blocks. MambaMixer [58] follows the core idea of MLP-Mixer [59] to combine Mamba-based token mixing with channel mixing, and follows the design of DenseNet [60] and DenseFormer [61] to allow blocks to access the previous features. Following [62], MambaR [63] evenly inserts register tokens into the input sequences and recycles registers for final decision predictions, which can help the model focus on more semantic regions. ARM [64] follows the GPT families [65, 66, 67], which utilizes the autoregressive modeling self-supervised framework. Similar to EViT [68], MambaPruning [69] introduces a token pruning method specifically designed for Mamba-based models. Inspired by [70], ShuffleMamba [71] introduces a training regularization technique that randomly shuffles token sequences during training to improve positional transformation invariance and overall model performance.
@@
On one hand, some papers focus on the micro-architecture designs to enhance the model’s ability. VSSD [72] introduces non-causal processing for image data by decoupling interaction magnitudes between hidden states and tokens. SparX-Mamba [73] introduces an efficient sparse cross-layer feature aggregation method specifically designed for SSMs. Vim-F [74] integrates frequency domain information into the basic block to achieve a better global receptive field. On the other hand, some work focuses on the macro-architecture advancement. SiMBA [75] utilizes hybrid sequence-channel modeling, effectively resolving the instability issues in large-size vision Mamba. StableMamba [76] establishes a robust interleaved Mamba-Attention framework that successfully scales SSMs to over 100 million parameters without resorting to knowledge distillation techniques. MambaVision [77] redesigns the Mamba architecture and integrates CNNs and ViT blocks on the architecture level, aiming at enhancing modeling capability.
@@ IV Micro-Architecture Advancement
This section focuses on the inner mechanism within the micro-architecture of Mamba block. Three core points are summarized: (1) SSM Formula Improvement (Section IV-A); (2) Scan Strategy (Section IV-B) and (3) Multi-Modal and Bi-Temporal Feature Interaction via Mamba (Section IV-C).
@@ IV-A SSM Formula Improvement
Figure 3: Overview of Recent Advancements in SSM Formulations. This figure categorizes improvements into four major aspects: (a) Context-SSM [78, 79] for local feature enhancement, integrating local context through CNNs; (b) Cross-SSM [80, 35, 36, 81, 82] for multi-modal and bi-temporal feature interaction, enabling cross-modal/temporal feature interaction at the parameter level; (c) New discretization methods for SSM [83], introducing second-order Runge-Kutta (RK2) for improved continuous-time approximation; and (d) Alternative hidden state definitions [84], utilizing a Minimum Search Tree (MST) to redefine hidden state transitions. These refinements collectively enhance the efficiency and expressiveness of SSMs in various applications.
@@ IV-A1 Context-SSM for Local Feature Enhancement
Context-SSM aims to naturally incorporate local context information into modeling long-range dependencies. Mamba-MOCO [78], building upon insights from MambaIRv2 [85], identifies two critical insights for improving SSMs in CV: (1) the output matrix $C$ in Eq. 7 functions as a ”Query” for local context information, which can address the inherent limitation of transforming 2D images into 1D sequences; (2) enhancing local context allows the model to better focus on local interactions while preserving its ability to model global dependencies. In response to these insights, Mamba-MOC [78] introduces the Context-SSM, as depicted in Fig. 3.a. Context-SSM incorporates two new parameters, $P$ and $D$, which are generated by multi-scale CNNs and a $1\times 1$ CNN. The parameters $P$ and $D$ capture local context information, ensuring that the model can effectively capture both local context and long-range dependencies for the input data. Similarly, AFA-Mamba [79] simply adds local context features $eg(L_{p})$ with the output equation together, which improves the ability to understand local context.
@@ IV-B Scan Strategy
TABLE I: This table summarizes a total of 44 different scan strategies utilized in the field of remote sensing. All acronyms are clearly illustrated in Fig. 5-11. The terms WB and BB denote ’Within Block’ and ’Between Block’, respectively, and are discussed in detail in Section IV-B1. Arrows in the direction column indicate the starting token and initial scan directions, which are further explained in Section IV-B3.
Scan Strategy Related Work
@@
Figure 4: Illustration of the scan strategy pipeline, comprising five key components, including feature preprocessing, scan sampling, scan directions, scan pattern, and feature post-processing. These five stages essentially transform a 2D feature map into multiple 1D sequences that conform to Mamba’s processing architecture. For clarity, an additional example is provided demonstrating the case without any preprocessing or postprocessing operations.
@@ IV-B1 Feature Preprocessing
Unlike Transformers [4, 170] processing tokens in parallel, Mamba performs sequential computation, meaning that the initial scan order and the starting token influence subsequent token interactions. This category aims to optimize spatial-spectral feature extraction through feature map rearrangement [122, 150, 113]. HSRMamba [122] introduces the Global Spectral Reordering Mechanism (GSRM), which rearranges spectral features based on the average correlation matrix values between spectrum, ensuring that highly correlated pixels are positioned closer together, as illustrated in Fig. 5.E. This improves long-range spectral feature modeling. VMambaSCI [150] proposes Spectral Exchange (SE), which reorders spectral features using a predefined sequence, as shown in Fig. 5.F. This enhances the long-range modeling of spectral features within the same spatial channel. DTAM [113] introduces Dynamic Token Argumentation (DTA) to enhance global feature learning. As depicted in Fig. 5.G, DTA consists of two routes: the left route activates object-related tokens while masking the others, and the right route activates contextual tokens while using random shuffling and Adaptive Instance Normalization (AdaIN) [171] for augmentation. This approach helps the model focus on meaningful object-related features while mitigating spectral variability caused by acquisition conditions.